# Beyond Binary: Towards Fine-Grained LLM-Generated Text Detection via Role Recognition and Involvement Measurement

### v1

@@ 3.2.3. Linguistic Feature Comparison
To systematically illustrate the differences in content across various LLM roles, we introduce seven linguistic feature metrics. Table 1 presents a comparison of these linguistic features across the four types of news content in HNDC. Average word and sentence count measure the number of words and sentences in a news article. Sentiment polarity score represents the emotional tone of a text, ranging from -1 to 1, with higher values indicating more positive sentiment, and lower values reflecting more negative sentiment. Grammatical errors measure the number of grammatical mistakes that occur per 1,000 words. Syntactic diversity measure the structural complexity by analyzing clause patterns. Vocabulary richness measures lexical diversity, ranging from 0 to 1, with higher values indicating greater lexical variation. Readability score measures the complexity of a text, with higher values indicating greater reading difficulty. The detailed calculation methods for the linguistic features are provided in Appendix C.
As shown in Table 1, we observe that LLM-generated news articles are generally shorter and contain fewer sentences compared to human-written news. In contrast, LLM-polished and LLM-extended news, incorporating more human inputs, are significantly richer and more comprehensive. The various types of news exhibit trivial differences in their sentiment polarity scores. From other linguistic features, human writing shows greater lexical and syntactic variation with lower reading difficulty, whereas LLM writing is more standardized, featuring fewer grammatical errors and minimal use of informal writing styles.
@@ 4. EXPERIMENTS
In this section, we evaluate the performance and generalization of 10 baseline detection methods (§4.1) in our LLMDect framework across the two proposed detection tasks. Firstly We train supervised detection methods on the HNDC and evaluate their test set performance, while also reporting the zero-shot LLM detector’s results (§4.2). Then we evaluate the generalization and robustness of the best-performing detection models on the HNDC across two dimensions in DetectEval: cross-context (§4.3) and multi-intensity(§4.4). Furthermore, we discuss the data leakage issue when using LLMs as zero-shot detectors (§4.5) to ensure fairness in detection outcomes. Finally, we evaluate their effectiveness by using them as feature extractors (§4.6). We report the F1 score for each LLM role and evaluate the performance of each detection method on the LLM-RR task using the weighted F1 score. We report the Mean Squared Error (MSE) and Mean Absolute Error (MAE) on the LLM-IM task.
4.1. Baseline Detection Methods
We consider 10 baseline detection methods. To illustrate the difficulty of the two proposed detection tasks, we first focus on decoder-only LLMs, which have demonstrated exceptional performance across a range of NLP tasks, including four recent advanced models Mistral (Jiang et al., 2023) (Mistral-7B-Instruct-v0.3), DeepSeek (DeepSeek-AI et al., 2024) (DeepSeek-V2-Chat as of June 28, 2024), LLaMa-3 (Dubey et al., 2024) (Meta-LLaMA-3-70B-Instruct), and GPT-4o (OpenAI, 2024) (as of May 05, 2024). We utilize these models in a zero-shot detection pattern using specifically designed prompts, as shown in Appendix D.
@@ 4.1. Baseline Detection Methods
cross-cultural German 96.22 94.87 98.18 0.0159 0.0135 0.0066
@@ 4.2. HNDC Performance Evaluation
Table 2 shows the detection performance of 10 baseline methods on the HNDC test set for the two tasks. The results show that supervised methods outperform the zero-shot LLM detector, with fine-tuned PLM-based models consistently achieving superior performance across both tasks. In contrast, advanced LLMs face challenges in accurately detecting content they have generated themselves. In the zero-shot LLM detector setting, Mistral and Deepseek show almost no detection capability, particularly compared to a base detector that predicts an LIR of 0 for all cases, indicating a total inability to detect LLM-generated content in the LLM-IM task. GPT-4o, while demonstrating limited ability to differentiate between human-authored and LLM-generated content, struggles to detect human-LLM collaboration. When providing detection rationales, GPT-4o tends to classify content as human-authored if it includes specific details, such as citations or data, while fluent and structured content is more likely to be identified as LLM-generated. Consequently, GPT-4o encounters significant challenges in detecting complex cases of human-LLM collaboration when relying on surface-level features alone. Feature-based models show intermediate performance, with varying detection capabilities across different LLM-generated content types. Notably, detectors using only linguistic features, without language models involved, still achieve objective results, indicating discernible linguistic differences between human-authored and LLM-assisted content, providing an interpretable basis for detection. In contrast, fine-tuned PLM-based detectors exhibit outstanding detection performance across all LLM roles, indicating that traditional methods remain highly effective in detecting LLM-generated content, even in the current era of advanced LLMs.
@@ 4.3. Cross-context Generalization Evaluation
We apply the three best-performing PLM-based detectors trained on HNDC to five distinct types of cross-context variations in DetectEval to assess their generalization capability. Table 3 presents the generalization evaluation results. Except for the cross-domain scenario, the PLM-based detectors demonstrated strong generalization across the other four variations, achieving 90% overall. Notably, the F1 score reached 95% in the cross-time, cross-prompt, and cross-source scenarios. The high performance in these scenarios suggests that PLM-based detectors are adaptable across different contexts. Specifically, in the cross-source scenario, detectors trained on corpora generated by weaker LLMs can effectively detect content from stronger LLMs, reducing computational resources while maintaining accuracy. Interestingly, in cross-cultural settings, we find that news from countries with higher visibility, such as Germany and China, is easier to identify, while news from Qatar presents more challenges. Besides, the lower performance in the cross-domain scenario likely results from differences in language structures and terminologies, highlighting the challenge of achieving robust cross-domain adaptability and the need for domain adaptation techniques to improve detection. By averaging the generalization performance of PLM-based detectors across various cross-context groups, we find that the DeBERTa-based detector exhibit the strongest generalization capability. We hypothesize that this may be attributed to DeBERTa’s use of relative position encoding, which improves its ability to capture long-range dependencies more effectively (He et al., 2020). Furthermore, we input the test data from HNDC into three original, non-fine-tuned PLMs to extract their inherent learned representations, which were then visualized using t-SNE after dimensionality reduction, as shown in Figure 2. Figure 2 shows that DeBERTa achieves clearer cluster separation compared to the other PLMs, indicating stronger discriminative ability. This suggests that DeBERTa captures relevant features more effectively, explaining its superior generalization in the cross-context evaluations.
@@ 4.6. LLMs Feature Extractors Analysis
Although generative decoder-only LLMs perform poorly in zero-shot detection, fine-tuning these models is computationally expensive, and PLM-based detection models have already achieved outstanding performance. Nevertheless, we can explore the potential of using these LLMs directly as feature extractors. Specifically, drawing on the GLTR (Gehrmann et al., 2019) approach, we train detection models using rank-based features extracted from different LLMs. The detection performance of each model when used as a feature extractor is presented in the Figure 4. We observe that decoder-only models, such as GPT-2, significantly outperform encoder-only models like RoBERTa, DeBERTa, and Longformer, even with comparable parameters [^11] https://huggingface.co/transformers/v4.11.3/pretrained_models.html. Among categories of LLM-assisted writings, LLM-creator texts are easiest to distinguish, whereas distinguishing between LLM-Polisher and Human-Author texts remains challenging. Additionally, comparisons among different size of GPT-2 reveal that larger models demonstrate better feature effectiveness.
@@ 5. Conclusion
In this paper, we introduce a new detection paradigm that moves beyond binary classification by considering both the role and level of LLM involvement in content creation. We proposed two novel tasks: LLM Role Recognition (LLM-RR) and LLM Influence Measurement (LLM-IM), offering a more fine-grained approach to detecting LLM-generated content. To support these tasks, we develop LLMDetect, a benchmark combining the Hybrid News Detection Corpus (HNDC) and DetectEval, designed to assess model robustness and generalization across diverse contexts. Our empirical evaluation of 10 baseline detection models demonstrated that fine-tuned PLM-based methods outperform others, with DeBERTa excelling in cross-context generalization and Longformer performing best with varying intensity levels. As LLM-generated content becomes more prevalent, particularly on social media, these findings highlight the importance of developing more effective and fine-grained detection models. Our approach provides valuable tools for detecting LLM involvement, contributing to improved content integrity in digital platforms.
Appendix A Prompts for HNDC Construction
@@
System Prompt: You are an AI assistant tasked with generating news articles. Given a news article title and its description, your task is to craft a well-structured and informative news article. Aim for a balanced and informative article that provides context and clarity to the reader. Adapt the tone and style to fit the nature of the news, whether it’s business, education, or scientific to engage the target audience effectively.
User Prompt: Here is a news article title: <title> and its description <description>, write a <category> news article based on this news article title and description I gave you and return news article as well as your title with the format Title: __ ### Article: __ (make sure to use ### as the delimiter). The article should reflect information available up to <publish date>.
@@
Prompt: Please rewrite the following sentences. ”’<news articles>”’
@@
System Prompt: You are an AI assistant tasked with generating news articles. Your task is to continue writing from the given incomplete news article and ensure the continuation is well-structured and informative. Aim for a balanced and informative article that provides context and clarity to the reader. Adapt the tone and style to fit the nature of the news, whether it’s business, education, or scientific to engage the target audience effectively.
User Prompt: Please complete the following news article. Don’t return the given text. The news begin with: ”’<beginning text>”’ Continue from here:
@@
The designed role-playing prompts for HNDC construction are shown in Table 5.
Appendix B Diverse Prompts
LLM-Creator
Prompt1: Here is a news article title: <title> and its description <description>, write a <category> news article based on this news article title and description I gave you and return news article as well as your title with the format Title: __ ### Article: __ (make sure to use ### as the delimiter). The article should reflect information available up to <publish date>.
Prompt2: Here is a news article title: <title> and its description <description>, write a <category> news article based on this news article title and description I gave you and return news article as well as your title with the format Title: __ ### Article: __ (make sure to use ### as the delimiter).
Prompt3: Please write a news about <title>, <description>. Return news article as well as your title with the format Title: __ ### Article: __ (make sure to use ### as the delimiter).
Prompt4: Here is a news article title: <title> and its description <description>, write a news article in an expert confident voice. Return news article as well as your title with the format Title: __ ### Article: __ (make sure to use ### as the delimiter).
Prompt5: Please write a news about <title>, <description> in a formal scientific writing voice. Return news article as well as your title with the format Title: __ ### Article: __ (make sure to use ### as the delimiter).
LLM-Polisher
Prompt1: Please polish the following sentences. ”’<news article text>”’
Prompt2: Please enhance fluency of the following sentences. ”’<news article text>”’
Prompt3: Please adjust structures of the following sentences. ”’<news article text>”’
Prompt4: Please rewrite the following sentences in a formal scientific writing voice. ”’<news article text>”’
Prompt5: Please polish the following sentences in a humorous voice. ”’<news article text>”’
LLM-Extender
Prompt1: Please complete the following news article. Don’t return the given text. The news begins with: ”’<beginning text>”’ Continue from here.
Prompt2: Please directly continue to write the news (not repeat my provided content): ”’<beginning text>”’
Prompt3: Please complete the following news article. Don’t return the given text. The text begin with: ”’<beginning text>”’
Prompt4: Complete the following unfinished news article. Don’t return the given text. The news begin with: ”’<beginning text>”’ Continue from here.
Prompt5: Please directly continue to write the news (not repeat my provided content): The news begin with: ”’<beginning text>”’ Continue from here.
Table 6. The Diverse Prompts for Cross-prompt of DetectEval
We assign five prompts to each role, with each news in the test set paired with a single prompt and call LLaMA-3 to generate the cross-prompt dataset. The prompts are shown in Table 6.
@@ Appendix C Supplement on Linguistic Features
Appendix D LLM Detectors Prompt
Table 7 presents the instruction prompts used by the zero-shot LLM detector for the two detection tasks.
@@ C. AI-Polished: The article is polished by AI from a human-written draft;
Please directly give the answer with answer-rationale pair in JSON format, with the structure: ”answer”: …, ”rationale”: ….”
Please directly give the ”answer” with ”A”, ”B”, ”C”, or ”D”, and explain your choice in two or three sentences (string format) in ”rationale”.
@@
The evaluation scores range from 0 to 1, where 0 indicates the article is completely human-written, and 1 indicates it is entirely AI-created.
Please directly give the answer with score in JSON format, with the structure: ”score”: …
@@
Appendix E GLTR Visualization

### v2

@@ 3.2.3. Linguistic Feature Comparison
To illustrate content differences across LLM roles, we introduce seven linguistic feature metrics. Table 1 compares these features across the four news types in HNDC. These metrics include average word and sentence count, sentiment polarity score (ranging from -1 to 1), grammatical errors per 1,000 words, syntactic diversity, vocabulary richness (0 to 1), and readability score. Calculation methods are detailed in Appendix C. As shown in Table 1, LLM-generated news is generally shorter with fewer sentences compared to human-written news. LLM-polished and LLM-extended news, with more human input, are richer and more comprehensive. Sentiment polarity scores show minimal differences, while human writing has greater lexical and syntactic variation with lower reading difficulty. LLM writing is more standardized, with fewer grammatical errors and less informal style.
@@ 4. EXPERIMENTS
cross-cultural Germany 96.22 94.87 98.18 0.0159 0.0135 0.0066
@@
In this section, we evaluate the performance and generalization of 10 baseline detection methods (§4.1) in our LLMDetect framework across the two proposed detection tasks. Firstly We train supervised detection methods on the HNDC and evaluate their test set performance, while also reporting the zero-shot LLM detector’s results (§4.2). Then we evaluate the generalization and robustness of the best-performing detection models on the HNDC across two dimensions in DetectEval: cross-context (§4.3) and multi-intensity(§4.4). Furthermore, we discuss the data leakage issue when using LLMs as zero-shot detectors (§4.5) to ensure fairness in detection outcomes. Finally, we evaluate their effectiveness by using them as feature extractors (§4.6). We report the F1 score for each LLM role and evaluate the performance of each detection method on the LLM-RR task using the weighted F1 score. We report the Mean Squared Error (MSE) and Mean Absolute Error (MAE) on the LLM-IM task.
4.1. Baseline Detection Methods
We consider 10 baseline detection methods. To illustrate the difficulty of the two proposed detection tasks, we first focus on decoder-only LLMs, which have demonstrated exceptional performance across a range of NLP tasks, including four recent advanced models Mistral (Jiang et al., 2023) (Mistral-7B-Instruct-v0.3), DeepSeek (DeepSeek-AI et al., 2024) (DeepSeek-V2-Chat as of June 28, 2024), LLaMa-3 (Dubey et al., 2024) (Meta-LLaMA-3-70B-Instruct), and GPT-4o (OpenAI, 2024) (as of May 05, 2024). We utilize these models in a zero-shot detection pattern using specifically designed prompts, as shown in Tabel 6.
@@ 4.2. HNDC Performance Evaluation
Table 2 shows the detection performance of 10 baseline methods on the HNDC test set for the two tasks. The results show that supervised methods outperform the zero-shot LLM detector, with fine-tuned PLM-based models consistently achieving superior performance across both tasks. In contrast, advanced LLMs face challenges in accurately detecting content they have generated themselves. In the zero-shot LLM detector setting, Mistral and Deepseek show almost no detection capability, particularly compared to a base detector that predicts an LIR of 0 for all cases, indicating a total inability to detect LLM-generated content in the LLM-IM task. GPT-4o, while demonstrating limited ability to differentiate between human-authored and LLM-generated content, struggles to detect human-LLM collaboration. When providing detection rationales, GPT-4o tends to classify content as human-authored if it includes specific details, such as citations or data, while fluent and structured content is more likely to be identified as LLM-generated. Consequently, GPT-4o encounters significant challenges in detecting complex cases of human-LLM collaboration when relying on surface-level features alone. Feature-based models show intermediate performance, with varying detection capabilities across different LLM-generated content types. Notably, detectors using only linguistic features, without language models involved, still achieve objective results, indicating discernible linguistic differences between human-authored and LLM-assisted content, providing an interpretable basis for detection. In contrast, fine-tuned PLM-based detectors exhibit outstanding detection performance across all LLM roles, indicating that traditional methods remain highly effective in detecting LLM-generated content, even in the current era of advanced LLMs. We hypothesize that the high F1 scores of these SOTA detectors result from our initial focus on longer news articles, which allow the model to capture more distinctive features, with performance declining as text length shortens (Wang et al., 2024).
[…]
@@ 4.3. Cross-context Generalization Evaluation
We apply the three best-performing PLM-based detectors trained on HNDC to five distinct types of cross-context variations in DetectEval to assess their generalization capability. Table 3 presents the generalization evaluation results. Except for the cross-domain scenario, the PLM-based detectors demonstrated strong generalization across the other four variations, achieving 90% overall. Notably, the F1 score reached 95% in the cross-time, cross-prompt, and cross-source scenarios. The high performance in these scenarios suggests that PLM-based detectors are adaptable across different contexts. Specifically, in the cross-source scenario, detectors trained on corpora generated by weaker LLMs can effectively detect content from stronger LLMs, reducing computational resources while maintaining accuracy. Interestingly, in cross-cultural settings, news from countries with higher visibility, like Germany and China, is easier to identify, while news from Qatar poses more challenges. This pattern is consistent when evaluating intersectional aspects. [^8] Detection performance on the test set, considering both out-of-domain LLMs and cultural aspects is shown in Appendix Table 7. Besides, the lower performance in the cross-domain scenario likely results from differences in language structures and terminologies, highlighting the challenge of achieving robust cross-domain adaptability and the need for domain adaptation techniques to improve detection. By averaging the generalization performance of PLM-based detectors across various cross-context groups, we find that the DeBERTa-based detector exhibit the strongest generalization capability. We hypothesize that this may be attributed to DeBERTa’s use of relative position encoding, which improves its ability to capture long-range dependencies more effectively (He et al., 2020). Furthermore, we input the test data from HNDC into three original, non-fine-tuned PLMs to extract their inherent learned representations, which were then visualized using t-SNE after dimensionality reduction, as shown in Figure 2. Figure 2 shows that DeBERTa achieves clearer cluster separation compared to the other PLMs, indicating stronger discriminative ability. This suggests that DeBERTa captures relevant features more effectively, explaining its superior generalization in the cross-context evaluations. [^9] Appendix Figure 6 shows the t-SNE visualization of the fine-tuned version, highlighting improved distinguishability.
[…]
@@ 4.6. LLMs Feature Extractors Analysis
Although generative decoder-only LLMs perform poorly in zero-shot detection, fine-tuning these models is computationally expensive, and PLM-based detection models have already achieved outstanding performance. Nevertheless, we can explore the potential of using these LLMs directly as feature extractors. Specifically, drawing on the GLTR (Gehrmann et al., 2019) approach, we train detection models using rank-based features extracted from different LLMs. The detection performance of each model when used as a feature extractor is presented in the Table 4. We observe that decoder-only models, such as GPT-2, significantly outperform encoder-only models like RoBERTa, DeBERTa, and Longformer, even with comparable parameters [^13] https://huggingface.co/transformers/v4.11.3/pretrained_models.html. Among categories of LLM-assisted writings, LLM-creator texts are easiest to distinguish, whereas distinguishing between LLM-Polisher and Human-Author texts remains challenging. Additionally, comparisons among different size of GPT-2 reveal that larger models demonstrate better feature effectiveness.
@@ 5. Conclusion
In this paper, we introduce a new detection paradigm that moves beyond binary classification by considering both the role and level of LLM involvement in content creation. We proposed two novel tasks: LLM Role Recognition (LLM-RR) and LLM Involvement Measurement (LLM-IM), offering a more fine-grained approach to detecting LLM-generated content. To support these tasks, we develop LLMDetect, a benchmark combining the Hybrid News Detection Corpus (HNDC) and DetectEval, designed to assess model robustness and generalization across diverse contexts. Our empirical evaluation of 10 baseline detection models demonstrated that fine-tuned PLM-based methods outperform others, with DeBERTa excelling in cross-context generalization and Longformer performing best with varying intensity levels. As LLM-generated content becomes more prevalent, particularly on social media, these findings highlight the importance of developing more effective and fine-grained detection models. Our approach provides valuable tools for detecting LLM involvement, contributing to improved content integrity in digital platforms.
Appendix A Detailed Explanations of Jaccard Distance
The Jaccard Distance, which is the complement of the Jaccard similarity, used to measure dissimilarity:
(5) $J\left(A,B\right)=1-\frac{\left|A\cap B\right|}{\left|A\cup B\right|}$
The value ranges from [0, 1]. A and B are more dissimilar as the value approaches 1, and more similar as it approaches 0. In our paper, we use the Jaccard Distance to quantify text modification before and after revision by comparing the word sets in the original and revised texts. Specifically, let $A$ and $B$ represent the sets of words in the human-written and LLM-polished documents, respectively.
Appendix B Implementation details
B.1. Prompts for Data Construction
@@
You are an AI assistant tasked with generating news articles. Given a news article title and its description, your task is to craft a well-structured and informative news article. Aim for a balanced and informative article that provides context and clarity to the reader. Adapt the tone and style to fit the nature of the news, whether it’s business, education, or scientific to engage the target audience effectively.
Here is a news article title: <title> and its description <description>, write a <category> news article based on this news article title and description I gave you and return news article as well as your title with the format Title: __ ### Article: __ (make sure to use ### as the delimiter). The article should reflect information available up to <publish date>.
@@
Please rewrite the following sentences.
”’<news articles>”’
@@
You are an AI assistant tasked with generating news articles. Your task is to continue writing from the given incomplete news article and ensure the continuation is well-structured and informative. Aim for a balanced and informative article that provides context and clarity to the reader. Adapt the tone and style to fit the nature of the news, whether it’s business, education, or scientific to engage the target audience effectively.
Please complete the following news article. Don’t return the given text. The news begin with: ”’<beginning text>”’
Continue from here:
@@
The role-playing prompts for HNDC construction are shown in Table 5. Table 6 shows the instruction prompts used by the zero-shot LLM detector.
B.2. Model setups
For LLM-generated news, we use LLama3-8b-Instruct on local A800 GPUs to generate data, with half-precision inference enabled. The maximum token output is set to 2048, and a termination symbol is specified. Additionally, sampling is enabled with the temperature of 0.6 and the top_p of 0.9. When fine-tuning PLM models, we use a batch size of 16, the Adam optimizer with a learning rate of 5e-5, and a dropout rate of 0.1. All models are fine-tuned for 5 epochs on A800 GPUs, with the best-performing model selected based on validation set performance.
@@ Appendix C Supplement on Linguistic Features
Figure 6. t-SNE visualization of representation from the fine-tuned DeBERTa.
[…]
@@ C. AI-Polished: The article is polished by AI from a human-written draft;
Please directly give the answer with answer-rationale pair in JSON format, with the structure: ”answer”: …, ”rationale”: ….” Please directly give the ”answer” with ”A”, ”B”, ”C”, or ”D”, and explain your choice in two or three sentences (string format) in ”rationale”.
@@
The evaluation scores range from 0 to 1, where 0 indicates the article is completely human-written, and 1 indicates it is entirely AI-created. Please directly give the answer with score in JSON format, with the structure: ”score”: …
@@
Culture Source LLM-RR (F1↑) LLM-IM (MSE↓)
RoBERTa DeBERTa Longformer RoBERTa DeBERTa Longformer
German GPT-4o 92.31 93.85 96.43 0.0152 0.0141 0.0078
Claude 93.07 92.32 95.89 0.0195 0.0198 0.0107
China GPT-4o 94.60 88.12 96.47 0.0134 0.0091 0.0057
Claude 96.16 88.27 96.62 0.0149 0.0121 0.0079
Qatar GPT-4o 84.91 88.15 74.13 0.0235 0.0161 0.0306
Claude 85.89 87.75 74.13 0.0285 0.0203 0.0331
Table 7. Detection performance on the test set, considering both out-of-domain LLMs and cultural aspects. Results show that out-of-domain cultural aspects are more challenging than LLMs, suggesting weaker LLMs can still be used for data generation.
Appendix D GLTR Visualization